OpenStudy (anonymous):

What is the simplified form of 4/y^2-5/3y?

OpenStudy (anonymous):

is it \[\frac{ 4 }{ y^{2} }-\frac{ 5 }{ 3y }\]?

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

\[\frac{ 4 }{ y^{2} }-\frac{ 5 }{ 3y }\] multiply first with 3 and second with y \[\frac{ 12 }{ 3y^{2} }-\frac{ 5y }{ 3y^{2} }\] --- \[\frac{ 12-5y }{ 3y^2 }\]

OpenStudy (anonymous):

Thank you, goodness, you're a life saver! :D

OpenStudy (anonymous):

uw ^^

OpenStudy (anonymous):

Great explanations! Really, thank you, I couldn't give thanks enough! I do an online school thing and they don't always provide the easiest to understand methods of solving